OpenStudy (anonymous):

8TH GRADE MATH... FIND 2 CONSECUTIVE ODD INTEGERS SUCH THAT 4 TIMES THE LARGER IS 29 MORE THAN 3 TIMES THE SMALLER. PLEASE EXPLAIN HOW YOU GOT YOUR ANSWER

hero (hero):

4(x+2) = 3(x) + 29

OpenStudy (anonymous):

CAN YOU EXPLAIN

hero (hero):

two consecutive odd integers such as 3,5 or 7,9 or 9,11.... notice that the second one is always two more than the first, so the general expression for consecutive integers is (x)+ (x+2), where for any set of two consecutive odd integers, x is the small integer, and x + 2 is the large integer.

hero (hero):

4 times the larger means we take the representation of the larger integer, x+2, and multiply it by 4. To multiply the whole expression by four, we have to put the whole expression in parentheses, then multiply it by four: (x+2)*4 or simply 4(x+2)

hero (hero):

Now, it says that the expression 4(x+2) is equal to three times the smaller plus 29. The smaller value is x. Remember, is means = So to transform the word problem into an equation, you have to follow the steps literally. If you do that, you'll end up with the expression I provided to you above.

hero (hero):

So to finish posting my solution: 4(x+2) = 3(x) + 29 4x + 8 = 3x + 29 x = 21 x + 2 = 23

hero (hero):

You can put the numerical value for x back into the original equation to see if it's true

OpenStudy (anonymous):

ok, thanks! :)

hero (hero):

4(21 + 2) = 3(21) + 29 4(23) = 63 + 29 92 = 92

hero (hero):

It's true, so we know we did it correctly

hero (hero):

By the way, math doesn't really have any problems. It does, but they're usually able to be solved using math.

OpenStudy (anonymous):

true..lol

OpenStudy (anonymous):

thanks once again